Derive Newton’ s second law in three dimension using the Lagrangian method.[6]


1
Expert's answer
2021-06-28T13:20:05-0400
"L=T-V\\\frac{\partial L}{\partial v}=mv\\\\\frac{\partial L}{\partial x}=-\frac{d U}{dx}\\m\frac{d v}{dt}=-\nabla V\\ma=F"

so we again arrive at Newton’s second law, F = ma, now in three dimensions.